OpenStudy (anonymous):

Simplify: Negative three and one-half minus negative twelve and one-fourth. Negative fifteen and three-fourths Fifteen and three-fourths Negative eight and three-fourths Eight and three-fourths

OpenStudy (anonymous):

Okay. First, change the mixed numbers into improper fractions. -3 and 1/2 = -7/2 -12 and 1/4 = -49/4 Now, your problem will read (-7/2) - (-49/4) You can change it to this: (-7/2) + (49/4) Next, you need to make a common denominator. Multiply the top and bottom of (-7/2) by 2. This will equal (-14/4). Now your problem says (-14/4) + (49/4) -14 + 49 = 35 35/4 = 8.75 or 8 and 3/4

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @FirelessDragon

OpenStudy (anonymous):

make it easier for u guys